x | y how do i solve this'[
3 |5.4
7 |12.6
12|21.6

Answer by checkley77(12844)

SEEING AS THIS IS A LINEAR RELATIONSHIP
5.4/3=1.8
12.6/7=1.8
21.6/12=1.8
WE CAN US THE LINE EQUATION:
Y=mX+b WHERE m=SLOPE & b=Y INTERCEPT.
SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(12.6-5.4)/7-3)
SLOPE=7.2/4
SLOPE=1.8
NOW WE SUBSTITUTE ONE SET OF X & Y VALUES & SOLVE FOR b.
5.4=1.8*3+b
5.4=5.4+B
b=0
THUS THE EQUATION IS:
Y=1.8X